car from rest is moving at a velocity of 86 km/hr in 10 seconds. Determine the acceleration of a car.

1
Expert's answer
2021-11-16T13:40:52-0500

Given:

"v_i=0"

"v_f=86\:\rm km/h=23.9\: m/s"

"t=10\:\rm s"


The acceleration of the car

"a=\frac{v_f-v_i}{t}=\frac{23.9-0}{10}=2.39\:\rm m/s^2"